Reports 1-1 of 1 Clear search Modify search
AdV-TCS (TCS control system)
berni, capocasa, gouaty, rocchi, spinicelli, taranto - 15:55 Wednesday 15 June 2022 (56150) Print this report
West arm matching measurement with HWS-INJ

In preparation for the afternoon shift, at 06.15 UTC all WI actuators were switched off.
The goal of the shift was to check the WI DAS setting for the compensation of the CH @ 93mW (needed to mimic the YAG at CARM NULL) with the HWS-INJ.
At the shift beginning, the HWS-INJ illumination was checked but no beam could be seen. So, Alessio and Piernicola went to EIB to fix the butterfly basis of the SLED beam and the beam reappeared on the sensor.
After that, there were some problems with the IR locking. Then the shift started later.
At 17.20 UTC a west arm scan with the green laser to measure the cold matching (used as a reference) was performed. 
The matching was 3.7% on Narm and 4% on warm  both values were higher than that measured on 6/07 (entry 56071)  (Narm: 2.5% Warm  1.3%) when the matching was recovered after PR RH switch off.
This is somehow expected for west arm as the matching was optimised with TCS on but it should not be the case for the north arm since TCS is in the same state as that used during matching optimisation.

The arms were locked on IR and drift control on the arms was activated (as seen in entry 55908).
At 17.31 UTC HWS-INJ acquisition started and at 17.35 UTC WI CH was switched on @ 93 mW. When the curvature reached the steady state, I switched on the WI DAS IN @ 400 mW and OUT @ 1100 mW.
At 20.49 UTC ITF went down and the HWS-INJ acquisition was stopped.
The entire behaviour of the curvature is shown in figure 1.

It seems to corresponds to what is measured by the matching (fig 2) and it shows that more DAS should be used to compensate the CH.

At 21.23 the initial TCS powers CH @ 120 mW, DAS IN @ 460 mW and OUT @ 1015 mW were set.

With the help of Romain Gouaty, we took the chance to check the beam on the cameras in transmission during TEM02 scan (Fig3:  B7_Cam2, Fig4: B8_Cam1). They show an astigmatic beam. 
B8 figure is less clear as, given to the different splitting ratio of the beam splitter on the end benches,  we have 10 time less power than on B7.

 

Images attached to this report
Comments to this report:
capocasa - 19:26 Wednesday 15 June 2022 (56161) Print this report

Today (15/05), I did a TEM02 scan to measure the maching with TCS, from 11.06 UTC to 12.02 UTC.

The measuremnet is in the attached plot. Narm: 3.8%. Warm: 2.1%.  TCS condition should be the same as the one of the measurment of 56071 where Narm was 2.5% and Warm 1.3%.

It is not clear what could have changed the matching from that time this but it looks quite a common effect.

 

 

Images attached to this comment
Search Help
×

Warning

×